# First-order quantum breakdown of superconductivity in amorphous superconductors

### v2

@@ Abstract
Continuous quantum phase transitions are widely assumed and frequently observed in various systems of quantum particles or spins [1, 2]. Their characteristic trait involves scaling laws governing a second-order, gradual suppression of the order parameter as the quantum critical point is approached. The localization of Cooper pairs in disordered superconductors and the resulting breakdown of superconductivity have long stood as a prototypical example [2, 3, 4, 5, 6]. Here, we show a departure from this paradigm, showcasing that amorphous superconducting films of indium oxide undergo a distinctive, discontinuous first-order quantum phase transition tuned by disorder. Through systematic measurements of the plasmon spectrum in superconducting microwave resonators, we provide evidence for a marked jump of both the zero-temperature superfluid stiffness and the transition temperature at the critical disorder. This discontinuous transition sheds light on the previously overlooked role of repulsive interactions between Cooper pairs and the subsequent competition between superconductivity and insulating Cooper-pair glass. Furthermore, our investigation shows that the critical temperature of the films no longer relates to the pairing amplitude but aligns with the superfluid stiffness [7], consistent with the pseudogap regime of preformed Cooper pairs [8, 9, 10, 11, 6]. Our findings raise fundamental new questions into the role of disorder in quantum phase transitions [12] and carry implications for superinductances in quantum circuits [13, 14, 15, 16, 17].
Superconductors undergo substantial changes in response to an increase in material disorder. Electron scattering, caused by disorder, increases resistivity and eventually leads to the breakdown of superconductivity due to Anderson localization and interactions [6]. This breakdown, commonly referred to as the superconductor-to-insulator transition, has long been considered a prototypical continuous quantum phase transition [2], tunable by disorder, magnetic field or charge carrier density [3, 4, 5, 6].
@@
In recent years, a body of work on thin films of various materials has revealed a more subtle interplay between phase and amplitude suppression [6]. Tunneling spectroscopy experiments have provided evidence of pairing amplitude persisting across the transition [20, 10, 21, 22], indicative of the localization of Cooper pairs in the insulator, together with strong spatial fluctuations [20, 10, 21, 23, 24, 25] and a pseudogap of preformed pairs[8, 9, 10, 26, 11, 25]. Concomitantly, this was accompanied by a substantial suppression of the superfluid (phase) stiffness [26, 27, 28, 25, 29] upon approaching critical disorder, pointing to a prevailing role of phase fluctuations. Yet, scaling laws, the hallmark of continuous QPTs, have never been demonstrated for the superfluid stiffness [30] in disorder-tuned transitions.
@@
At the core of this study is the systematic and accurate measurement of the superfluid stiffness, $\Theta$, together with the DC transport properties. We designed superconducting microwave stripline resonators (see Fig. 1b) of a:InO enabling us to directly extract the kinetic inductance per square of the materials, $L_{K}$, through the superconducting plasmon dispersion [31, 32]. In the two-dimensional limit, the superfluid stiffness straightforwardly follows from
@@
($\hbar$ the reduced Planck constant, $e$ the electron charge). Comparison of $\Theta$ with the other relevant energy scales of the superconducting state, such as the pairing gap and the superconducting transition temperature provides a direct assessment of the role of phase fluctuations. $\Theta$ smaller than $\Delta$ signals a phase-driven superconducting transition and the presence of preformed pairs [7]. It is worth noting that the effective dimensionality depends on the physical phenomenon under consideration: in our samples, it is 3D for Cooper pairing, 2D for phase fluctuations and 1D for plasmons (see Methods for discussion).
@@
Figure 2: Phase-driven superconducting transition. Experimental phase diagram representing the critical temperature $T_{c}$, the single-particle tunneling gap $\Delta$ (from [10, 22, 34]) and the low temperature superfluid stiffness $\Theta$ as a function of sheet resistance. Upon increase of disorder, that is, resistance, the superconductor goes from a BCS regime with $\Theta>\Delta$ and $T_{c}\propto\Delta$ to a phase fluctuation regime characterized by $\Theta<\Delta$ and $T_{c}\approx\Theta$. In the latter, the superconducting transition is driven by the establishment of phase stiffness of preformed Cooper pairs that emerge as a pseudogap in the single particle density of states. Blue, green and red dashed lines are guide for the eyes. Inset: tunneling conductance measured on an a:InO film ($T_{\rm c}=1.7$ K) as a function of bias voltage across the tunneling junction and temperature. The black dashedline marks the tunneling spectrum at $T=T_{c}$. The pseudogap above $T_{\rm c}$ extends up to $\Delta\simeq 6$ K in excellent agreement with the phase diagram. Dotted lines below (blue) and above $T_{c}$ (red) on the inset are also reported in the figure at the corresponding disorder. Figure reproduced from Ref. [10].
@@
where $\Delta_{c}$ is the superconducting collective gap right before the transition to insulator, which is significantly different from the single-particle spectral gap [11]. Here, $\nu_{0}$ is the density of states at the Fermi level, and $\varepsilon$ is the macroscopic dielectric constant in the insulating state. Note that Eq. (2) takes into account the fact that the elementary charge in our insulator is $2e$.
@@
A first-order transition is also anticipated in other disordered materials exhibiting an anomalously long electric screening length $l_{scr}$ comparable to the superconducting coherence length. In a:InO, calculations in Refs. [49, 51], along with estimates for the dielectric constant $\varepsilon$ in the insulating state[46, 53], suggest $l_{scr}\approx\xi_{loc}$, where the localization length $\xi_{loc}$ is approximately $4-5$ nm, close to that of the superconducting coherence length [34]. Recent experiments in granular aluminum films also reported such a nm-scale screening length [56]. This material shares similar features with a:InO, such as a discontinuous drop in $T_{c}$ at the superconductor-insulator transition and non-monotonous re-entrant insulating state [57], suggesting a first-order breakdown of superconductivity, here too. Interestingly, $\mathrm{LaAlO_{3}/SrTiO_{3}}$ heterostructures [28] reach 80 nH per square at the transition, i.e. $\Theta\sim 0.1$ K, which is consistent with a smaller Coulomb gap due to the large dielectric constant of SrTiO3 known to be $\varepsilon\approx 10^{4}$.
@@ Discussion
In a strongly disordered low-dimensional system, a first-order transition is not expected as disorder is known to smear the transition due to disorder-induced energy fluctuations and the pinning of domain walls that separate competing phases [58, 59]. However, in a three-dimensional (3D) system as ours, these effects do not necessarily preclude the existence of a sharp first-order transition in the thermodynamic sense. Yet, observation of other distinct signatures of the first-order phase transition described above, such as glassy dynamics arising from the very slow motion of domain walls, would require the ability to continuously tune parameters such as magnetic field [60] or carrier density via a gate voltage [28], which is beyond the scope of this study.
An important theoretical insight from our findings is the significant role of disorder in quantum phase transitions, which exceeds the expectations set by the standard mapping between $T=0$ quantum statistics of a $D$-dimensional system and the classical statistical mechanics of the same system in $D+1$ dimensions [1]. This is because frozen disorder is time-independent, leading to world-line correlations in $D+1$ dimensional space-time. In contrast, pure point-like-correlated disorder in $D+1$-dimensional space-time of an auxiliary classical system lacks such persistent correlations in the time domain. This makes disorder markedly more influential in altering the nature of a quantum phase transition compared to a clean system, as observed in temperature-driven classical transitions.
Finally, the emergence of materials with very large kinetic inductance, resulting from suppressed superfluid stiffness, underscores their potential significance for both quantum circuits and sub-THz photon detectors. Superinductors capable of maintaining a quality factor of $\gtrsim 10^{4}$ (see Extended Data Figure 4 and Methods for a:InO quality factor), combined with their compact footprint, offer versatility across various applications, ranging from inductively shunted qubits to dissipative resonators and nonlinear parametric amplification, as well as highly sensitive photon detectors. Our findings establish an upper limit for the highest achievable kinetic inductance in disordered materials. Nevertheless, the precise role of disorder in bulk dissipation remains a critical aspect yet to be fully understood [61], promising to stimulate further investigation and reveal new fundamental insights.
@@ Methods
$K_{nm}=3\alpha\left(1-\frac{1}{4}\delta_{nm}\right)\frac{\xi^{2}}{Lw}\,\frac{\hbar\omega_{n}\omega_{m}}{\Theta},$ (5)
where $\xi$ is the dirty limit superconducting coherence length, $\Theta$ is the superfluid stiffness, $\alpha=\frac{\pi}{4}+\frac{3}{4\pi}\approx 1.02$ is determined by the current-phase nonlinearity of a diffusive superconductor, and $\delta_{nm}=\begin{cases}0\text{ if }n\neq m\\ 1\text{ if }n=m\end{cases}$.
For a strongly disordered superconductor, the semiclassical description is rendered inapplicable both due to strong localization and due to the presence of pseudogap, hence the exact value of $K_{nm}$ is not known. However, one can still use Eq. (5) to estimate the order of magnitude of the effect.
@@
The model above accounts accurately for the geometry of our samples, leaving the kinetic inductance per square $L_{K}=lw$ as the only fitting parameter. The reliability of extraction of $L_{K}$ has been confirmed by electromagnetic simulations in this geometry, see Supplementary Information I. We also note that the obtained $L_{K}$ are orders of magnitude larger than geometric inductances.
@@ Acknowledgments
We thank D. Basko, L. Benfatto, J. Delahaye, T. Grenet, V. Kravtsov, M. Muller, M. Scheffler, and C. Strunk for valuable discussion. We thank E. Eyraud for assistance with cryogenics and J. P. Martinez for preliminary measurements. T.C. and B.S. acknowledge funding from the ANR Project No. ANR-19-CE30-0014-CP-Insulators. A.V.K. is grateful for the support by Laboratoire d’excellence LANEF in Grenoble (ANR-10-LABX-51-01). B.S. has received funding from the European Union’s Horizon 2020 research and innovation program under the ERC grant SUPERGRAPH No. 866365. N.R. has received funding from the European Union’s Horizon 2020 research and innovation program under the ERC grant SuperProtected No. 101001310. N.R., B.S. and D.P. acknowledge funding from the ANR agency under the ’France 2030’ plan, with Reference No. ANR-22-PETQ-0003.
@@
.1 Electromagnetic simulation of plasmon dispersion relations
@@
.2 Extraction of quality factors
@@
.3 Model for quality factors as function of temperature and power
The model used to fit the quality factors shown in Extended Fig. 4a and b a and b consists of three dissipation channels added in parallel:
@@
.4 Dissipation in a:InO resonators is not due to surface dielectric loss
The microscopic origin of TLSs in our resontaors is not known. Studies show that dielectric materials located at the interfaces surrounding the resonators host such defects and contribute to microwave dissipation [67]. These dielectric materials come in thin layers (typically a few nm-thick), and originate from native oxides at the substrate or metal surfaces, or from resist residues. Here we show that the anomalously large dissipation in indium oxide resonators cannot be explained by surface dielectric loss alone.

### v3

@@ Abstract
Continuous quantum phase transitions are widely assumed and frequently observed in various systems of quantum particles or spins. Their characteristic trait involves scaling laws governing a second-order, gradual suppression of the order parameter as the quantum critical point is approached. The localization of Cooper pairs in disordered superconductors and the resulting breakdown of superconductivity have long stood as a prototypical example. Here, we show a departure from this paradigm, showcasing that amorphous superconducting films of indium oxide undergo a distinctive, discontinuous first-order quantum phase transition tuned by disorder. Through systematic measurements of the plasmon spectrum in superconducting microwave resonators, we provide evidence for a marked jump of both the zero-temperature superfluid stiffness and the transition temperature at the critical disorder. This discontinuous transition sheds light on the previously overlooked role of repulsive interactions between Cooper pairs and the subsequent competition between superconductivity and insulating Cooper-pair glass. Furthermore, our investigation shows that the critical temperature of the films no longer relates to the pairing amplitude but aligns with the superfluid stiffness, consistent with the pseudogap regime of preformed Cooper pairs. Our findings raise fundamental new questions into the role of disorder in quantum phase transitions and carry implications for superinductances in quantum circuits.
Superconductors undergo substantial changes in response to an increase in material disorder. Electron scattering, caused by disorder, increases resistivity and eventually leads to the breakdown of superconductivity due to Anderson localization and interactions [1]. This breakdown, commonly referred to as the superconductor-to-insulator transition, has long been considered a prototypical continuous quantum phase transition [2, 3], tunable by disorder, magnetic field or charge carrier density [2, 4, 5, 6, 1].
@@
In recent years, a body of work on thin films of various materials has revealed a more subtle interplay between phase and amplitude suppression [1]. Tunneling spectroscopy experiments have provided evidence of pairing amplitude persisting across the transition [9, 10, 11, 12], indicative of the localization of Cooper pairs in the insulator, together with strong spatial fluctuations [9, 10, 11, 13, 14, 15] and a pseudogap of preformed pairs [16, 17, 10, 18, 19, 15]. Concomitantly, this was accompanied by a substantial suppression of the superfluid (phase) stiffness [18, 20, 21, 15, 22] upon approaching critical disorder, pointing to a prevailing role of phase fluctuations [7]. Yet, scaling laws, the hallmark of continuous QPTs, have never been demonstrated for the superfluid stiffness [23] in disorder-tuned transitions.
@@
[…]
We designed superconducting microwave stripline resonators (see Fig. 1b) of a:InO enabling us to directly extract the kinetic inductance per square of the materials, $L_{K}$, through the superconducting plasmon dispersion [25, 26]. In the two-dimensional limit, the superfluid stiffness straightforwardly follows from
@@
($\hbar$ the reduced Planck constant, $e$ the electron charge).
@@
Figure 2: Phase-driven superconducting transition. Experimental phase diagram representing the critical temperature $T_{c}$, the single-particle tunneling gap $\Delta$ (from [10, 12, 27]) and the low temperature superfluid stiffness $\Theta$ as a function of sheet resistance. The tunneling data were obtained on virtually identical films [10, 12, 27] allowing a consistent comparison with the current microwave samples. Upon increase of disorder, that is, resistance, the superconductor goes from a BCS regime with $\Theta>\Delta$ and $T_{c}\propto\Delta$ to a phase fluctuation regime characterized by $\Theta<\Delta$ and $T_{c}\approx\Theta$. In the latter, the superconducting transition is driven by the establishment of phase stiffness of preformed Cooper pairs that emerge as a pseudogap in the single particle density of states. Blue, green and red dashed lines are guide for the eyes. Inset: tunneling conductance measured on an a:InO film ($T_{\rm c}=1.7$ K) as a function of bias voltage across the tunneling junction and temperature. The black dashedline marks the tunneling spectrum at $T=T_{c}$. The pseudogap above $T_{\rm c}$ extends up to $\Delta\simeq 6$ K in excellent agreement with the phase diagram. Dotted lines below (blue) and above $T_{c}$ (red) on the inset are also reported in the figure at the corresponding disorder. Figure reproduced from Ref. [10].
@@
where $\Delta_{c}$ is the superconducting collective gap right before the transition to insulator, which is significantly different from the single-particle spectral gap [19]. Here, $\nu_{0}$ is the density of states at the Fermi level, and $\varepsilon$ is the macroscopic dielectric constant in the insulating state. Note that Eq. (2) takes into account the fact that the elementary charge in our insulator is $2e$. It is important to highlight that the long-range Coulomb interaction driving the first-order transition is the interaction between bound electron pairs. The presence of a pseudogap in our strongly disordered films (see inset of Fig. 2 and references [10, 19]) provides experimental evidence for the existence of these pairs. Theoretically, Ref. [45] explains why short-range electron repulsion [8] doesn’t destroy the pairing itself. In short, this resilience is due to the fractal nature of electron wavefunctions in near-critical Anderson insulators.
@@
A key implication of our work is that the properties of the disorder-tuned SIT depend on microscopic details of the material under consideration, and therefore are not universal. For instance, for a first-order transition to occur, two conditions must be met: the superconductor must present a pseudogap (like it is the case for NbN, TiN, grAl and some others), and the Coulomb gap amplitude must be comparable to the superfluid stiffness (as stated by Eq. (3)). These two conditions strongly depend on parameters of the material such as dielectric constant and density of states. A first-order transition is anticipated in disordered materials exhibiting an anomalously long electric screening length $l_{scr}$ comparable to the superconducting coherence length. In a:InO, calculations in Refs. [43, 45], along with estimates for the dielectric constant $\varepsilon$ in the insulating state [40, 47], suggest $l_{scr}\approx\xi_{loc}$, where the localization length $\xi_{loc}$ is approximately $4-5$ nm, close to that of the superconducting coherence length [27]. Recent experiments in granular aluminum films also reported such a nm-scale screening length [50]. This material shares similar features with a:InO, such as a discontinuous drop in $T_{c}$ at the superconductor-insulator transition and non-monotonous re-entrant insulating state [51], suggesting a first-order breakdown of superconductivity, here too. Interestingly, $\mathrm{LaAlO_{3}/SrTiO_{3}}$ heterostructures [21] reach 80 nH per square at the transition, i.e. $\Theta\sim 0.1$ K, which is consistent with a smaller Coulomb gap due to the large dielectric constant of SrTiO3 known to be $\varepsilon\approx 10^{4}$.
@@ Discussion
In a strongly disordered low-dimensional system, a first-order transition is not expected as disorder is known to smear the transition due to disorder-induced energy fluctuations and the pinning of domain walls that separate competing phases [52, 53, 54]. However, in a three-dimensional (3D) system as ours, these effects do not necessarily preclude the existence of a sharp first-order transition in the thermodynamic sense. Yet, observation of other distinct signatures of the first-order phase transition described above, such as glassy dynamics arising from the very slow motion of domain walls, would require the ability to continuously tune parameters such as magnetic field [55] or carrier density via a gate voltage [21], which is beyond the scope of this study.
An important theoretical insight from our findings is the significant role of disorder in quantum phase transitions. Standard descriptions involve a mapping of QPTs in the D-dimensional system at T=0 onto the classical thermal transition in dimension D + 1 (where the ”1” denotes the time dimension) [2, 3]. Such correspondence is possible in clean systems because world-lines in D+1 dimensional space-time are not correlated in the time domain. In contrast, strong, ”frozen” disorder is by definition time-independent, leading to world-line correlations in D + 1 dimensional space-time. This makes disorder markedly more influential in altering the nature of a quantum phase transition compared to a clean system and thus may call to revisit our understanding of disorder-tuned QPTs.
Finally, the emergence of materials with very large kinetic inductance, resulting from suppressed superfluid stiffness, underscores their potential significance for both quantum circuits and sub-THz photon detectors [56, 57, 58, 59, 60]. Superinductors capable of maintaining a quality factor of $\gtrsim 10^{4}$ (see Extended Data Figure 4 and Methods for a:InO quality factor), combined with their compact footprint, offer versatility across various applications, ranging from inductively shunted qubits to dissipative resonators and nonlinear parametric amplification, as well as highly sensitive photon detectors. Our findings establish an upper limit for the highest achievable kinetic inductance in disordered materials. Nevertheless, the precise role of disorder in bulk dissipation remains a critical aspect yet to be fully understood [61], promising to stimulate further investigation and reveal new fundamental insights.
@@ Methods
The technique employs a microwave resonator with a simple straight-strip line geometry, which can be probed at very low temperatures (20 mK), very low excitation powers (down to a few photons in average) and GHz frequencies (quantum regime $\hbar\omega\gg k_{B}T$). The cryostat in which the measurements took place is routinely used for the characterization of high-coherence superconducting circuits, providing a suitable filtered microwave environment. The originality of our experimental technique is the exploitation of the intrinsic non-linearities of the superconductor (Kerr effect) using two-tone measurements for the detection of resonance modes over a large frequency range, far beyond the frequency bandwidth allowed by standard microwave components.
@@
$K_{nm}=3\gamma\left(1-\frac{1}{4}\delta_{nm}\right)\frac{\xi^{2}}{Lw}\,\frac{\hbar\omega_{n}\omega_{m}}{\Theta},$ (5)
where $\xi$ is the dirty limit superconducting coherence length, $\Theta$ is the superfluid stiffness, $\gamma=\frac{\pi}{4}+\frac{3}{4\pi}\approx 1.02$ is determined by the current-phase nonlinearity of a diffusive superconductor, and $\delta_{nm}=\begin{cases}0\text{ if }n\neq m\\ 1\text{ if }n=m\end{cases}$.
In this estimation we have assumed that the kinetic inductance fraction $\alpha=L_{K}/L_{\text{tot}}=1$, i.e. the geometric inductance of the wire is negligible compared to the kinetic inductance, as is the case for disordered indium oxide. For a strongly disordered superconductor, the semiclassical description is rendered inapplicable both due to strong localization and due to the presence of pseudogap, hence the exact value of $K_{nm}$ is not known. However, one can still use Eq. (5) to estimate the order of magnitude of the effect.
@@
The model above accounts accurately for the geometry of our samples, leaving the kinetic inductance per square $L_{K}=lw$ as the only fitting parameter. The reliability of extraction of $L_{K}$ has been confirmed by electromagnetic simulations in this geometry (see Supplementary Information I) and therefore offers high accuracy in determining the superfluid stiffness. This accuracy is further demonstrated by the excellent agreement between microwave and DC measurements. Note that, contrary to most other experimental probes of the superfluid response (two-coil mutual inductance, scanning SQUID, microwave spectrometer…), we do not need any calibration of the experimental setup or background subtraction in order to extract the kinetic inductance. We also note that the obtained $L_{K}$ are orders of magnitude larger than geometric inductances, hence confirming the assumption $\alpha=1$.
@@
To summarize, one should consider our system to be effectively 1D, 2D or 3D depending on the mechanism at play. While long-wavelength plasmons responsible for microwave resonances and the Kerr effect are one-dimensional, the effect of thermal phase fluctuations near $T_{c}$ giving rise to the relation $T_{c}=\Theta$ is of a 2D nature, as expected from the BKT mechanism. Microscopic effects of disorder, local superconducting pairing and Coulomb repulsion occur on shorter scales than the film dimensions and are 3D in nature.
@@ Acknowledgments
We thank D. Basko, L. Benfatto, J. Delahaye, T. Grenet, V. Kravtsov, M. Muller, M. Scheffler, and C. Strunk for valuable discussion. We thank E. Eyraud for assistance with cryogenics and J. P. Martinez for preliminary measurements. T.C. and B.S. acknowledge funding from the ANR Project No. ANR-19-CE30-0014-CP-Insulators. A.V.K. is grateful for the support by Laboratoire d’excellence LANEF in Grenoble (ANR-10-LABX-51-01). B.S. has received funding from the European Union’s Horizon 2020 research and innovation program under the ERC grant SUPERGRAPH No. 866365. N.R. has received funding from the European Union’s Horizon 2020 research and innovation program under the ERC grant SuperProtected No. 101001310. N.R., B.S. and D.P. acknowledge funding from the ANR agency under the ’France 2030’ plan, with Reference No. ANR-22-PETQ-0003. I.P. acknowledges support by the Deutsche Forschungsgemeinschaft (DFG) via the grant MI 658/14-1.
@@
I Electromagnetic simulation of plasmon dispersion relations
@@
II Supplementary two-tones spectroscopy data
SI Data Fig. 2: Two-tones spectroscopy spectrum for four samples of increasing disorders. From top to bottom: resonance modes obtained via the amplitude or phase response of two-tones spectroscopy for increasingly disordered samples TC001, TC007, DP-res11 and TC016-6. Data for TC016-6 is filtered and a background is removed.
Figure 2 shows additional two-tone spectroscopy data for four other samples with increased disorder (TC001, TC007, DP-res11, and TC016-6), complementing Figure 1c which displayed data for sample TC014. The resistances and geometries of these samples are provided in Extended Data Table 1. Either the phase or the amplitude of the transmission coefficient $S_{21}$ is represented, depending on which parameter exhibited the best signal-to-noise ratio. Whether a dip or a peak is observed in the transmission depends on the frequency of the probe tone (i.e., whether it sits exactly at the resonance frequency or is slightly detuned).
For a given geometry, the plasmon spectrum is given by $2\pi f_{n}=vk_{n}$ where $k_{n}=n\pi/L$ is the wavevector for mode $n$ and $v=1/\sqrt{lc_{k}}$. Therefore, at low $k$, the mode spacing is roughly constant $\sim v\pi/L$, which is close to the fundamental mode frequency $2\pi f_{1}=v\pi/L$. At a given wire width, the fundamental frequency and mode spacing are controlled by two parameters: the kinetic inductance per unit length $l$ and the resonator length $L$. As disorder increases, $l$ increases, which reduces the frequency. However, our resonators do not all have the same length (see Extended Data Table 1). In particular, our most strongly disordered resonator just before the SIT (sample TC016-6) is shorter than most others, which allows for a mode spacing large enough to resolve multiple resonances.
III Extraction of quality factors
@@
IV Model for quality factors as function of temperature and power
The model used to fit the quality factors shown in Extended Fig. 4a and b consists of three dissipation channels added in parallel:
@@ IV.2 Two-Level Systems
V Dissipation in a:InO resonators is not due to surface dielectric loss
The microscopic origin of TLSs in our resonators is not known. Studies show that dielectric materials located at the interfaces surrounding the resonators host such defects and contribute to microwave dissipation [67]. These dielectric materials come in thin layers (typically a few nm-thick), and originate from native oxides at the substrate or metal surfaces, or from resist residues. Here we show that the anomalously large dissipation in indium oxide resonators cannot be explained by surface dielectric loss alone.